OpenStudy (anonymous):

One number is 4 less than 3 times a second number. If 3 more than twice the first number is decreased by the second, the result is 11. Use the subsitution method. What Is The Number

OpenStudy (anonymous):

Im So Confused lol

OpenStudy (anonymous):

Here are The Choices 3 4 8 11

jimthompson5910 (jim_thompson5910):

let x and y be the two numbers

OpenStudy (anonymous):

Ok

jimthompson5910 (jim_thompson5910):

One number is 4 less than 3 times a second number means that x = 3y - 4

OpenStudy (anonymous):

ok Yes Im Following

jimthompson5910 (jim_thompson5910):

oh wait, made a typo, one sec

jimthompson5910 (jim_thompson5910):

3 more than twice the first number is decreased by the second, the result is 11 translates to 2x + 3 - y = 11 now you plug in x = 3y - 4 to get 2x + 3 - y = 11 2(3y - 4) + 3 - y = 11 6y - 8 + 3 - y = 11 5y - 3 = 11 ... y = ??? tell me what you get

OpenStudy (anonymous):

3

jimthompson5910 (jim_thompson5910):

oh sry, made another typo and it should be 5y - 5 = 11 but I'm not getting a whole number when I solve for y, one sec

jimthompson5910 (jim_thompson5910):

maybe it's more like this

jimthompson5910 (jim_thompson5910):

3 more than twice the first number is decreased by the second, the result is 11 translates to 2(x-y) + 3 = 11

OpenStudy (anonymous):

Okay

jimthompson5910 (jim_thompson5910):

if so, then 2(x-y) + 3 = 11 2(3y - 4-y) + 3 = 11 2(2y - 4) + 3 = 11 4y - 8 + 3 = 11 4y - 5 = 11 4y = 16 y = 4 x = 3y - 4 x = 3*4 - 4 x = 8 so it looks like the first number is 8

OpenStudy (anonymous):

Okay Thanks

jimthompson5910 (jim_thompson5910):

yw